Brief Exercise 14-3 Your answer is incorrect. Try again. Monty Corporation issues $450,000 of 1396 bonds that are due in 9 years and pay interest semi-annually. At the time of issue, the market rate for such bonds is 12%. Click here to view the factor table. Using time value of money tables, a financial calculator, and computer spreadsheet functions, calculate the bonds issue price. (For calculation purposes, use 5 decimal places as displayed in the factor table provided and round final answer to 0 decimal places, e.g. 5,275.) Issue price of bond 270571
Answer #1

The issue price of the bond is calculated as sum of present value of all the cash flows receivable from bond till maturity.

The discounting factor that will be used to calculate the present value is the market rate of such bonds, i.e., 12%. Since the interest is payable semi-annually the discounting rate used to calculate the present value is 6% (half of 12%).

Discounting factor is calculated using formula = 1/(1+i)^n

where i = rate of interest (6%)

n = number of periods (Since the interest payable semi-annually, the number of periods are 9*2 = 18)

Coupon payment semi-annually = $450,000*6.5% = $29,250

The issue price of the bond will be calculated as below: Period Cash DF @ 6% Present value flow 29250 29250 29250 29250 29250 29250 29250 29250 29250 29250 29250 29250 29250 29250 29250 29250 29250 27,594.4.5 26,032.50 24,558.89 23,168.63 21,857.36 20,620.08 19,453.01 18,351.74 17,313.08 16,332.91 15,408.61 14,536.37 13,713.57 12,937.28 12,205.15 11,514.26 10,862.28 0.350341,67,900.45 4,74,361 0.94340 0.89000 0.83962 0.79209 0.74726 0.70496 0.66506 0.62741 0.59190 0.55839 0.52679 0.49697 0.46884 0.44230 0.41727 0.39365 0.37136 1 2 3 4 6 8 10 12 13 15 16 17 18 479250 Issue price of the bond

For extra understanding:

When the coupon rate of the bond is more than the market interest rate, the bond will be issued at premium same like this example.

When the coupon rate is less than the market interest rate, the bond will be issued at discount to compensate the investor from less coupon payment.